# Back from the dead: AT2019aalc as a candidate repeating TDE in an AGN

### v1

@@ 4.5.3 Over-flare-2 spectra
We performed optical spectroscopic observations of AT2019aalc during its second flare. These are characterized by a blue continuum with several lines appearing in comparison to the spectra discussed above. We compare the second LRIS spectrum of AT2019aalc (covering the broadest wavelength range) with a composite SDSS spectrum based on Seyfert 1 galaxies published in Pol & Wadadekar (2017). This comparison shows similarities but also significant differences discussed below.
The line widths and strengths of the earlier detected broad Balmer emission lines (H, H and H) increased and both the line widths and strengths of the Balmer lines peaked days before the optical peak. Two lines of the series (H and H) were detected only in the over-flare-2 spectra.
Over the second optical flare of AT2019aalc we detected another BF line N iii together with the transition line He ii . This doublet has not been detected in the archival SDSS spectrum taken years before the discovery of the transient or in the post-flare-I spectrum. We detected this doublet already in the first spectrum taken during the second flare. The strengths of these lines increased significantly around the second optical flare and peaked on 2023-07-11, very close to the optical continuum peak (2023-07-13). As we detected the Bowen line N iii even days after the optical peak of the second flare, we can conclude the long-term persistence of at least one BF line after the second flare as well. In addition, the second LRIS spectrum shows a clear increase in the BF line O iii and in He ii when comparing it to the post-flare-1 spectrum. Both lines increased by a factor of with respect to the post-flare-1 LRIS spectrum. We might detected O iii which is also known as a Bowen line (Selvelli et al. 2007), however, due to a possible blending from Fe vii, its detection is questionable. Nevertheless, this Bowen line is created through the same channel (O1) as the above discussed O iii which suggests its detection. O iii was detected in some metal-rich TDEs before (Leloudas et al. 2019) such as AT2019dsg (Cannizzaro et al. 2021).
We detect the high-ionization coronal lines seen in the post-flare-1 LRIS spectrum over the second flare in each spectrum. Furthermore, a new coronal line (Fe xi, ionization potential: eV) appears on 2023-08-17, around a month after the peak of the second optical flare. Interestingly, the line intensities of these lines are larger after the optical peak than around the peak in contrast with the BF lines and the above-mentioned ionized helium lines which experienced their maximums very close to the continuum peak. Moreover, [Fe xi and [Fe x] are not just persistent but the strongest in the late-time DBSP spectrum, taken almost a year after the optical peak of the second flare. We detect the high-ionization coronal line [S xii only in the late-time spectrum. This line is an extreme ionization line with an ionization potential of eV thus implies that unusually high level of ionization is taken place in AT2019aalc almost a year after its second optical peak. For this coronal line we estimate FWHM(S xii) km s-1.
We detected the [Ne V] emission lines at and in the second LRIS spectrum, similarly to the post-flare-1 spectrum and implying continuous ionization of the EELR. The [O iii] line appears in the late-time spectrum.
@@ 5.1.1 Balmer lines
The line width of the Balmer lines and its limited variability are not common for regular TDEs and flaring AGN which tend to have decreasing and increasing Balmer line widths over time, respectively (Gezari 2021). However, this behavior is more compatible with the spectroscopic properties of the TDE-AGN PS16dtm (Blanchard et al. 2017). Still, the slight increasing and broadening indicate high-velocity gas moving in the BLR, close to the SMBH. In addition, a more complex structure of these lines seen in the late-time DBSP spectrum.
5.1.2 Bowen lines
@@
The Bowen Fluorescence mechanism in astrophysical environments is a good indicator of absorbed EUVsoft X-ray flux (at wavelengths shorter than the He ii Lyman limit of eV) as these high-energy photons are converted into the He ii Lyman emission required for the excitation of the O iii and N iii (Selvelli et al. 2007). This is consistent with the multiple soft X-ray flares and the extreme UV luminosities of AT2019aalc. The BF line O iii has been observed in the Bowen Fluorescence Flares AT2017bgt (Trakhtenbrot et al. 2019) and AT2021loi and only in a few Seyfert galaxies before (e.g., Malkan 1986; Schachter et al. 1990), however, we note that due to the atmospheric limit at these lines are not easy to detect. This line is typically much weaker in normal Seyfert galaxies as in TDEs, BFFs or in our case (see Fig. 14). The BF line N iii is not generally seen in AGN (e.g., Vanden Berk et al. 2001), however, has been observed in several optical TDEs (Charalampopoulos et al. 2022b) and other SMBH-related transients; the BFFs AT2017bgt, F01004-2237, OGLE17aaj (Trakhtenbrot et al. 2019) and AT2021loi (Makrygianni et al. 2023), the peculiar transient event AT2019avd (Malyali et al. 2021) and a few flaring NLSy1s (Frederick et al. 2021). In a TDE case, the disrupted star’s material forms an accretion disk and the photosphere absorbs the EUVsoft X-ray photons originating from accretion processes and this way provide a plausible explanation for the appearance of the BF lines. BFFs have been first identified by Trakhtenbrot et al. (2019) who revealed three flares (AT2017bgt, F01004-2237 and OGLE17aaj) in AGN with the appearance of persistent Bowen lines. Later, Makrygianni et al. (2023) studied a transient, AT2021loi, classified as a BFF. The detection of a BF line almost years after the first optical flare of AT2019aalc suggests its long persistence, similar to the cases of the BFFs studied by Trakhtenbrot et al. (2019). In the case of AT2017bgt, this Bowen line was still detected days after the discovery of the transient. These lines in BFFs identified by Trakhtenbrot et al. (2019) may originate from a pre-existing BLR (a region with dense line-emitting gas) which was suddenly exposed to the intense ionizing UV emission, which also indicates enhanced accretion onto the SMBH. Alternatively, the BF lines might be related to a newly launched outflow driven by a sudden increase in accretion rate which interacts with the BLR. We compare the over-flare-2 high-resolution LRIS spectrum of AT2019aalc with the spectra of the BFFs AT2017bgt and AT2021loi, the BFF candidate AT2019avd, the TDE-Bowen AT2018dyb and the repeating partial TDE AT2020vdq in Fig. 15 with focusing on the N iii He ii emission line doublet. We further compare the second LRIS spectrum of AT2019aalc with the LRIS spectrum of AT2021loi in Fig. 16 zooming into the O iii and He ii region.
The increasing of these lines around the optical peak indicate that these lines were enhanced by the Bowen Fluorescence mechanism. The BF process is particularly strong in some He ii, O iii and N iii transitions. These lines are obviously detected during the second optical flare, while atypical for normal Seyfert 1 galaxies (see a comparison plot in Fig. 14). Moreover, the O iii lines at and normally detected in AGN do not show up in our spectra, similarly to the case of AT2017bgt. These results suggest that the BF mechanism plays an important role and this mechanism traces significantly enhanced accretion onto the SMBH suggesting the unusual nuclear transient nature of AT2019aalc.
@@
The He ii ionized helium line is presented in our spectra taken between the optical flares, and is also increased during the second optical flare. This emission line is commonly present in AGN and known as a recombination line, tracing ionized gas in the vicinity of the central SMBH responsible for the photoionization (e.g. Wang & Kron 2020). This line is known to be created by photoionization due to (soft) X-ray photons with a rough correspondence of He ii photon emitted for each keV X-ray photon (Pakull & Angebault 1986; Schaerer et al. 2019; Cannizzaro et al. 2021). In the spectra of AT2019aalc, He ii is present together with strong N iii and O iii BF lines and the less commonly detected He ii transition, indicating its relation to the Bowen Fluorescence mechanism. Furthermore, we calculated a maximum line intensity ratio of whose value significantly exceeds the typical ratios estimated for AGN (earlier Vanden Berk et al. (2001) estimated ratios of while Shirazi & Brinchmann (2012) gives ). Our value is, however, more comparable with the line ratio estimated for the BF transient AT2017bgt which is (Trakhtenbrot et al. 2019). The line intensities of He ii and N iii peak days before the continuum peak, suggesting extreme ionization by UV/soft X-ray photons around the peak. The He ii /He i ratio is sensitive to the change in the temperature and density as these lines belong to the same element but in two different ionization states (e.g., Ilić et al. 2009). We found the highest ratio close to the continuum peak, implying increased temperature in the BLR. The smallest ratio can be estimated from the post-flare-1 and late-time spectrum taken days after the continuum peak that is due to a drop in the temperature.
@@ 5.1.3 Coronal lines
We detected several high-ionization coronal lines. Two of the detected coronal lines present already in the post-flare-1 spectrum taken years after the first optical peak indicating a connection to the first optical flare.
@@
Based on a study of SDSS spectra, Wang et al. (2012) reported the discovery of Extreme Coronal Line Emitter (ECLE) galaxies which show extremely strong coronal lines. Three TDEs (AT2017gge, AT2019qiz and AT2022upj) were also found to have coronal lines (Onori et al. 2022; Newsome et al. 2022; Short et al. 2023). The BFF AT2021loi (Makrygianni et al. 2023) and the spectacular transient event AT2019avd (Malyali et al. 2021) have significantly detected coronal lines, too. Based on the publicly available spectra of AT2017bgt on the Transient Name Server111111https://www.wis-tns.org/ (TNS) we see enhanced coronal line emission in this BFF as well. Wang et al. (2012) states that the most possible origin of the ECLEs are remnants of an earlier TDE outburst which was later supported by Short et al. (2023) who revealed that the TDE AT2019qiz more likely follows the line ratio correlations of ECLEs instead of regular AGN. Interestingly, the coronal lines in AT2019qiz appeared around days after the optical peak, further supporting this scenario. Moreover, the estimated rate of coronal line emitters is consistent with the estimated rate of TDEs (Wang et al. 2012). According to Wang et al. (2012), the ECLEs could be explained through the coronal lines arising from Fe liberated from dust grains that were destroyed by the TDE flare. The large dust echo of AT2019aalc is consistent with this picture. Furthermore, the TDEs with coronal line detection clearly have more luminous IR dust echo compared to other optical TDEs. AT2019qiz is one of the few TDEs included in the accretion flare sample of van Velzen et al. (2024) due to its strong dust echo. Another coronal line-detected TDE AT2017gge is part of the mid-infrared (MIR) dust echo flare sample of Hinkle (2024) which contains ambiguous nuclear transients with high dust covering factors. The extreme dust echo seems to characterize the BFFs as well. The BFFs AT2021loi, AT2017bgt, OGLE17aaj and the candidate BFF AT2019avd are present in the above-mentioned MIR dust echo sample (Hinkle 2024) and of them are also part of the Flaires sample of dust-echo-like IR flares (Necker et al. 2024). The Bowen line transient AT2019pev that shows strong high-ionization coronal lines (Frederick et al. 2021; Yu et al. 2022) is also part of the MIR flare sample of Hinkle (2024). These findings suggest a connection not only between the ECLEs and TDEs, but also the BFFs. Fig. 17 shows the [Fe x] versus the [O iii luminosities of the EECLEs studied by Wang et al. (2012), AT2019qiz (Short et al. 2023), two BFFs and AT2019aalc together with the SDSS Seyfert sample of Gelbord et al. (2009). This clearly implies an offset of the ECLEs, BFFs and AT2019aalc from the regular AGN sample.
The [Fe x] line evolution is shown in Fig. 18. The strength of this coronal line in AT2019aalc is further increased during the second flare and seem to be peaking lately with respect to the continuum peak. The coronal line TDE AT2019qiz shows a late-time increase in the this line as well (Short et al. 2023). Similarly to the case of AT2019qiz, we also found that [Fe x line is the strongest high-ionization coronal line presented in the spectra, while [Fe vii is the weakest one. The coronal line [Fe x] has been detected in several AGN before, however, the maximum line intensity ratio we calculated ([Fe x / [O iii] ) clearly overshoots the maximum ratio (, Nagao et al. (2000) found for AGN. For AT2021loi, Makrygianni et al. (2023) calculated even higher ratios, up to .The line width of [Fe x] shows a clear evolution during the monitoring of AT2019aalc as we estimated values of FWHM([Fe x]. Interestingly, the line widths we estimated for [Fe x are clearly larger than the maximum values found for coronal line emitters (around km s-1, Wang et al. 2012), however, very consistent with the width estimated for AT2021loi. For AT2021loi, Makrygianni et al. (2023) gives a line width of km s-1 which implies that the line is consistent with a BLR instead of NLR origin since the latter would require the typical values estimated for coronal line emitters ( km s-1, Wang et al. 2012). The BLR origin would naturally explain the increasing emission in these lines after the optical peak (due to the light travel time) but is also consistent with the distance between the dust responsible for the reprocessed IR emission and the SMBH ( pc, van Velzen et al. 2024). This suggests a connection between the unusually large dust echos and the extreme coronal lines in these transient sources. An alternative explanation is collisional excitation of these coronal lines. This is compatible with a detailed multi-wavelength study of AT2019qiz which suggests that outflows (powered by accretion or stream collision) are responsible for driving the optical flare of this coronal-line TDE (Nicholl et al. 2020).
The sulfur line [S xii has the second highest ionization potential among the optical lines detected in AGN before (the highest one is [Ar xiv which we do not detect in AT2019aalc). We zoom into the SPS fitted DBSP spectrum to show the detection of this line in Fig. 19. We can reasonably assume that the most recent and very strong X-ray flare, started less than a month before the appearance of the sulfur line, is responsible for the creation of this emission line. Apart from ECLEs in the sample of Wang et al. (2012), [S xii] has only been reported in Seyfert galaxies (Kraemer & Crenshaw 2000; Mazzalay et al. 2010; Cerqueira-Campos et al. 2021; Oh et al. 2022) including the neutrino emitter (IceCube Collaboration et al. 2022) NGC 1068 and in the nuclear transients VT J1548 (Somalwar et al. 2022) and AT2022fpx (Koljonen et al. 2024). VT J1548 is a radio-detected transient that was classified as a promising BFF candidate and notably with remarkable TDE-like characteristics (Somalwar et al. 2022). Similarly, AT2022fpx shares properties both with TDEs and BFFs (Koljonen et al. 2024). To our knowledge, AT2019aalc is so far only the third nuclear transient detected with a line requiring such a high creation ionization potential.
5.1.4 Forbidden lines
We find evidence for the forbidden line [O iii] in the post-flare-1 LRIS spectrum and the late-time DBSP spectrum. Its detection, however, remains ambiguous because the Balmer line series are complex in our spectra, including blending from H. Nevertheless, the detections of [O iii] years after the first flare and year after the second flare is not unexpected. After all, this emission line requires densities much lower than the UV-line emitting clouds, and consistent with the outer NLR (or simple ISM) clouds (Ji et al. 2024) but is located inner than the emitting region of [O iii ] (Nagao et al. 2001). The line ratio of [O iii] /H suggests extremely hard ionizing radiation from the ionizing source. Using James Webb Space Telescope (JWST) observations, this line is sometimes detected and studied in detail in high- AGN (used to distinguish AGN from star-forming regions), however, such a ratio was detected only in a very few cases (Übler et al. 2024; Ji et al. 2024; Mazzolari et al. 2024). The [O iii ]/[O iii] ratio of is clearly higher than calculated for AGN which are typically below (Baskin & Laor 2005; Übler et al. 2024; Binette et al. 2024)).
The highly ionized neon lines typically originate from star formation activity or photoionization of the EELR by the AGN. In some extreme cases, high-velocity AGN-driven shocks that ionize extended emission regions are expected to be responsible for the presence of the lines [Ne V] at and (Maddox 2018). The strengths of these lines in AT2019aalc are higher than the [O ii] [O ii] doublets which indicate their origin related to AGN activity since [Ne V] is not excited by star formation (Maddox 2018). These shocks are possibly, but not exclusively, related to radio jets as high shock velocities of km s-1 are required to produce these lines (Maddox 2018). We estimated slightly lower velocities, which are therefore more consistent with the EELR origin and significant AGN photoionization in this outer region. From the increasing line intensities of [O iii] and [Ne V] we can conclude that the NLR and EELR are also strongly ionized, however, in the late-time spectrum these decreased faster than the Bowen line N iii. The ionization level of the BLR is therefore more persistent than in the outer regions.
@@ 5.2 Optical re-brightening
The long-term ZTF light curve of AT2019aalc shows two well separated flares. Both flares evolved quickly and reached their peaks within months, followed by a slow decay on yearly timescales. The first flare decayed for more than a year, while the second flare has been decaying for around a year.
Several ZTF monitored extragalactic transients have been found with bumps seen in their optical light curves (see Soraisam et al. (2022) for a recent review)121212We note that ZTF has not been running long enough to discover flare repeats within longer timescales e.g., a decade that actually might also be common. Most of them are SNe, however, a couple of them have been identified as TDEs (AT2020acka, AT2020nov and AT2021ehb). The latter two TDEs show prominent bumps in their light curves while decaying, while AT2020acka exhibits a long and monotonic bump. These TDEs are, however, experienced these bumps within a year from their peak. Moreover, the second flare of AT2019aalc peaked even at higher luminosities than the initial one. AT2020nov and AT2020acka experienced less significant second bumps, while AT2021ehb has two comparable peaks. The optical emission of these transients are poorly understood, and Soraisam et al. (2022) also states that their identifications as TDEs are ambiguous. Generally, TDEs are characterized by smoothly and rapidly declining light curves ( as expected from mass fallback considerations (Rees 1988) without any prominent bumps in the case of full disruption. AT2019aalc in turn has a slowly decaying first flare with a power-law fit returning a power-law index of and a second flare that has started years after the initial flare. The first systemically identified repeating partial TDE (AT2020vdq), however, shows not just a double-peaked long-term optical light curve, but its second flare is times more luminous compared to the first one (Somalwar et al. 2023). This implies an even more significant optical re-brightening than the one experienced by AT2019aalc. The second flares of partial TDEs are expected to follow an extremely steep power-law when declining. In the case of AT2020vdq, the second flare declined much slower (Somalwar et al. 2023) than this but still clearly faster than the second flare of AT2019aalc which decayed with a power-law index of . Other candidate repeating TDEs have been found with periodic flares at different wavelengths; ASASSN 14ko with more than observed optical/UV/X-ray flares (Payne et al. 2021), eRASSt J045650.3-20375 with X-ray/UV flares (Liu et al. 2023), AT2022dbl with two optical flares (Lin et al. 2024), while AT2018fyk (Wevers et al. 2019a, 2023) and RX J133157.6-324319.7 (Hampel et al. 2022) showed a one-time X-ray/UV and X-ray re-brightening episodes, respectively. In these cases, due to its grazing orbit, the star orbiting in the vicinity of the SMBH is thought to be only partially disrupted. After another pericenter passage, the surviving core thus feeds a new accretion flare, resulting in X-ray/UV re-brightening. For the Bowen-line and coronal-line emitter transient AT2019avd with two distinct optical flares, Chen et al. (2022) discusses a partial TDE scenario.
@@
Interestingly, most of the sources of the small sample of classified BFFs show optical re-brightening or bumps. Fig. 20 shows the optical evolution of the three BFFs with detected re-brightening episodes together with AT2019aalc. The optical light curve of AT2017bgt shows a bump roughly days after its initial flare which is also clearly seen in the binned light curve published in Trakhtenbrot et al. (2019). One of the two other BFFs originally studied by (Trakhtenbrot et al. 2019), F01004-2237, experienced a significant optical re-brightening around years after its initial flare detected in . Originally, Catalina Sky Survey (CSS) light-curve data revealed that this ULIRG had undergone a luminous continuum flare (Tadhunter et al. 2021) and later Trakhtenbrot et al. (2019) classified it as a BFF similar to AT2017bgt. The binned light curve of this source published in Makrygianni et al. (2023) shows a bump around a year after its peak. We investigated the recent activity of the source and found that it significantly re-brightened in based on its ATLAS monitoring. The recent flare was more luminous than the first one and decayed faster. In addition to these two BFFs, AT2021loi shows a bump during its decaying, roughly a year after the first flare reached its peak. Makrygianni et al. (2023) suggests that optical re-brightening or bumps seen when decaying might be common features of BFFs. Only one transient classified as a BFF, OGLE17aaj, has not shown signs of optical re-brightening or bumps in its light curve to date. The first flares of the studied BFFs decay on yearly timescales, i.e. much slower than TDE flares but more similar to AT2019aalc. However, the BFFs with re-brightening episodes tend to have a faster decaying second flare/bump in comparison with their initial flare, while in the case of AT2019aalc the second flare decays slower. The BFFs classified so far tend to have remarkable variances in their optical light curve evolution in addition to the similarities we see. These might be explained by different subclasses of BFFs, similarly to regular TDEs, where different subclasses have been defined based on their light curve evolution (e.g., Charalampopoulos et al. 2023).
5.3 Possible explanations of the long-term radio flare

### v2

@@ 4.5.3 Over-flare-2 spectra
We performed optical spectroscopic observations of AT2019aalc during the second flare. These are characterized by a blue continuum with several emission lines appearing in comparison to the spectra discussed above. We compare the second LRIS spectrum of AT2019aalc (covering the broadest wavelength range) with a composite SDSS spectrum based on Seyfert 1 galaxies published in Pol & Wadadekar (2017) ,in Fig. 10. This comparison shows similarities but also significant differences discussed below.
Over the second optical flare, strong Balmer lines were detected. Two lines of the series (H and H) were detected only in the over-flare-2 spectra.
Over the second optical flare of AT2019aalc, we detected the Bowen Fluorescence line complex N iii +He ii . This complex was not detected in the earlier spectra. The relative strength of the BF complex to H reached a maximum on 2023-07-11, very close to the optical continuum peak (2023-07-13). As we detected the Bowen line N iii even days after the optical peak of the second flare, we conclude that it persists over long timescales. We also find hints of O iii which is also known as a Bowen line (Selvelli et al. 2007), however, due to a possible blending from Fe vii, its detection is questionable. Nevertheless, this Bowen line is created through the same channel (O1) as the above discussed O iii which supports the hint of detection. O iii was detected in some metal-rich TDEs before (Leloudas et al. 2019) such as AT2019dsg (Cannizzaro et al. 2021).
We detect the high-ionization coronal lines seen in the post-flare-1 LRIS spectrum over the second flare in each spectrum. Furthermore, a new coronal line appears: (Fe xi (ionization potential: eV). We also detected the [Ne V] emission lines at and in the second LRIS spectrum, as well as in the post-flare-1 spectrum.
@@ 5.1.1 Balmer lines
AT2019aalc exhibits prominent Balmer lines in each spectrum. In addition, a more complex structure of these lines appears in the late-time over-flare-2 spectrum. The Balmer series lines here exhibit a redward wing which, however, in some cases, might result from blending with other lines, such as H and the BF line N,iii . Interestingly, the post-flare-I spectrum shows similar structure of the Balmer line series.
5.1.2 Bowen fluorescence lines
@@
Wavelength coincidences between emission lines (”line fluorescence”) can be an important source of radiative excitation. The Bowen Fluorescence mechanism in astrophysical environments is a good indicator of absorbed EUVsoft X-ray flux (at wavelengths shorter than the He ii Lyman limit of eV) as these high-energy photons are converted into the He ii Lyman emission required for the excitation of the O iii and N iii lines in the optical and NUV regimes (Netzer et al. 1985; Netzer 1990). This is consistent with the multiple soft X-ray flares and the extreme UV luminosities of AT2019aalc. The BF line O iii has been observed in the Bowen Fluorescence Flares AT2017bgt (Trakhtenbrot et al. 2019) and AT2021loi and only in a few Seyfert galaxies before (e.g., Malkan 1986; Schachter et al. 1990). However, we note that due to the atmospheric cutoff at these lines are difficult to detect. This line is typically weaker in regular Seyfert galaxies than in the case of AT2019aalc (see Fig. 10). The BF line N iii is not generally seen in AGN (e.g., Vanden Berk et al. 2001), however, it has been observed in several optical TDEs (Charalampopoulos et al. 2022b) and other SMBH-related transients: BFFs (Trakhtenbrot et al. 2019; Makrygianni et al. 2023), the peculiar transient event AT2019avd (Malyali et al. 2021) and a few flaring NLSy1s (Frederick et al. 2021).
The detection of a BF line almost years after the first optical flare of AT2019aalc suggests its long persistence, similar to the BFFs studied by Trakhtenbrot et al. (2019). In the case of AT2017bgt, this line was still detected days after the discovery of the transient. The Bowen features in F01004-2237 were detected even years after the initial flare (Tadhunter et al. 2021). We compare the over-flare-2 high-resolution LRIS spectrum of AT2019aalc to the spectra of BFFs, a TDE-Bowen, and a repeating partial TDE in Fig. 11 around the N iii He ii complex. We further compare the second LRIS spectrum of AT2019aalc with the LRIS spectrum of AT2021loi in Fig. 12 zooming into the O iii He ii region.
@@
The He ii ionized helium line is presented in the spectrum taken between the optical flares.. This emission line is commonly present in AGN and known as a recombination line, tracing ionized gas in the vicinity of the central SMBH responsible for the photoionization (e.g. Wang & Kron 2020). This line is known to be created by photoionization due to (soft) X-ray photons with a rough correspondence of He ii photon emitted for each keV X-ray photon (Pakull & Angebault 1986; Schaerer et al. 2019; Cannizzaro et al. 2021). In the spectra of AT2019aalc, He ii is detected together with strong N iii and O iii BF lines and He ii , indicating its relation to the Bowen Fluorescence mechanism. Furthermore, we calculated a maximum line intensity ratio of which significantly exceeds the typical ratios estimated for AGN (earlier Vanden Berk et al. (2001) estimated ratios of while Shirazi & Brinchmann (2012) gives ). Our value is, however, more comparable with the line ratio estimated for the BFF AT2017bgt (, Trakhtenbrot et al. 2019).
@@ 5.1.3 Coronal lines
Based on a study of SDSS spectra, Wang et al. (2012) reported the discovery of Extreme Coronal Line Emitter (ECLE) galaxies which show extremely strong coronal lines. Three TDEs (AT2017gge, AT2019qiz and AT2022upj) were also found to exhibit coronal lines (Onori et al. 2022; Newsome et al. 2022; Short et al. 2023). The BFF AT2021loi (Makrygianni et al. 2023) and the spectacular transient event AT2019avd (Malyali et al. 2021) also exhibit strong coronal lines. The BFF F01004-2237 showed variable coronal lines on a timescale of years. Based on the publicly available spectra of AT2017bgt on the Transient Name Server777https://www.wis-tns.org/ (TNS) we see enhanced coronal line emission in this BFF as well.
@@
Wang et al. (2012) state that the most possible origin of the ECLEs are remnants of an earlier TDE outburst which was later confirmed by Short et al. (2023) who revealed that the TDE AT2019qiz more likely follows the line ratio correlations of ECLEs instead of regular AGN. The coronal lines in AT2019qiz appeared around days after the optical peak, further supporting this scenario. Moreover, the estimated rate of coronal line emitters is consistent with the estimated rate of TDEs (Wang et al. 2012). Later, Callow et al. (2024) discovered 9 new ECLEs using SDSS spectra and refined the previous rate calculations. They found that the ECLE rates are below the TDE rates, however, can still be consistent with assuming that only to per cent of all TDEs produce variable coronal lines. According to Wang et al. (2012), the origin of the coronal lines in ECLEs can be explained with the liberation of Fe from dust grains that were destroyed by the TDE flare. A follow-up spectroscopic analysis of the original ECLE sample of Wang et al. (2012) revealed that the coronal line signatures did not recur in five out of the seven objects, supporting their transient light echo origin (Clark et al. 2024). The large dust echo of AT2019aalc is consistent with this picture. Furthermore, the TDEs with coronal line detection have more luminous IR dust echo compared to other optical TDEs. AT2019qiz is one of the few TDEs included in the accretion flare sample of van Velzen et al. (2024) due to its strong dust echo. Another coronal line-detected TDE AT2017gge is part of the mid-infrared (MIR) dust echo flare sample of Hinkle (2024) which contains ambiguous nuclear transients with high dust covering factors. The extreme dust echo seems to characterize the BFFs as well. The BFFs AT2021loi, AT2017bgt, OGLE17aaj and the candidate BFF AT2019avd are present in the above-mentioned MIR dust echo sample (Hinkle 2024) and of them are also part of the Flaires sample of dust-echo-like IR flares (Necker et al. 2024). The Bowen line transient AT2019pev shows strong high-ionization coronal lines (Frederick et al. 2021; Yu et al. 2022) and is also part of the MIR flare sample of Hinkle (2024). These findings suggest a connection not only between the ECLEs and TDEs, but also the BFFs. We performed modeling of the publicly available LRIS spectrum of AT2017bgt taken on 2017-02-23 using PyQSOFit, following the procedure described in Subsect. 3.5. Fig. 13 shows the [Fe x] versus the [O iii luminosities of the ECLEs studied by Wang et al. (2012), AT2019qiz (Short et al. 2023), two BFFs and AT2019aalc together with the SDSS Seyfert sample of Gelbord et al. (2009). This clearly implies an offset of the ECLEs, BFFs and AT2019aalc from the regular AGN sample.
@@ 5.2 Optical re-brightening
The long-term ZTF light curve of AT2019aalc shows two well-separated flares. Both flares evolved quickly and reached their peaks within months, followed by a slow decay on yearly timescales. Generally, TDEs are characterized by smoothly and rapidly declining light curves () as expected from mass fallback considerations (Rees 1988) without any prominent bumps in the case of full disruption. AT2019aalc in turn has a slowly decaying first flare with a power-law fit returning a power-law index of and a second flare that began years after the initial flare. The first systematically identified repeating partial TDE (AT2020vdq), however, shows not just a double-peaked long-term optical light curve, but its second flare is times more luminous compared to the first one (Somalwar et al. 2023). This implies an even more significant optical re-brightening than the one detected in AT2019aalc. The second flare of partial TDEs is expected to follow a steep declining with a power-law of . In the case of AT2020vdq, the second flare declined much slower (Somalwar et al. 2023) than this but still clearly faster than the second flare of AT2019aalc which declined with a power-law index of .
@@
Interestingly, most of the sources of the small sample of classified BFFs show optical re-brightening or bumps. Fig. 14 shows the optical evolution of the three BFFs with detected re-brightening episodes together with AT2019aalc. The optical light curve of AT2017bgt shows a bump roughly days after its initial flare which is also clearly visible in the binned light curve published in Trakhtenbrot et al. (2019). Originally, F01004-2237 had undergone a luminous flare (Tadhunter et al. 2021) in 2010 and later Trakhtenbrot et al. (2019) classified it as a BFF. The binned light curve of F01004-2237 published in Makrygianni et al. (2023) shows a bump around a year after its peak. We investigated the recent activity of this source and found that it significantly re-brightened in based on its ATLAS monitoring. The recent flare was more luminous than the first one and declined faster. In addition to these two BFFs, AT2021loi shows a bump when declining, roughly a year after the first flare reached its peak. Makrygianni et al. (2023) suggests that optical re-brightening or bumps seen when decaying might be common features of BFFs. Only one transient classified as a BFF, OGLE17aaj, has not shown signs of optical re-brightening or bumps in its light curve to date. The first flare of the studied BFFs decay on yearly timescales, much slower than regular TDE flares but more similar to AT2019aalc. However, the BFFs with re-brightening episodes tend to have a faster decaying second flare/bump in comparison with their initial flare, while in the case of AT2019aalc the second flare decays slower. The BFFs classified so far tend to have remarkable variances in their optical light curve evolution despite the similarities we see. These might be explained by different subclasses of BFFs, similarly to regular TDEs, where different subclasses were defined based on their light curve evolution (e.g., Charalampopoulos et al. 2023).
5.3 Radio characteristics
5.3.1 Possible explanations of the long-term radio flare